OpenStudy (anonymous):

Find the first three derivatives of f(x)=tan3x.

OpenStudy (mathmale):

To help you get started: look up the derivative of the tangent function. What is the derivative of the simpler function y = tan x? Now return to y = tan 3x. Since the argument (3x) is now a function in its own right, we need to apply the Chain Rule: Find the derivative of y = tan (3x) with respect to (3x). Then, multiply this by the derivative of the (separate) function (3x). Try this, please.

OpenStudy (anonymous):

Well I get the first derivative which is, 3sec^2(3x)

OpenStudy (anonymous):

I don't get how to do the second derivative.

OpenStudy (anonymous):

second derivative \[\frac{ 18\tan 3x }{ \cos ^{2}3x}\]

OpenStudy (mathmale):

Think of y' as y' = 3*(sec(3x))^2. This is a POWER FUNCTION. Apply the constant multiplier rule, the power rule and the chain rule, in that order, to determine the derivative of 3*(sec(3x))^2; the result will be y''.

OpenStudy (mathmale):

If ceyhun is correct, then y''' is found using the QUOTIENT rule (and several others).

OpenStudy (anonymous):

Well the answer in the book comes out to be, 18sec^2(3x)tan(3x)

OpenStudy (mathmale):

What's more important is how we arrive at that answer. Please consider how you'd do this using the basic rules of differentiation. Given the practice problem y = tan x / cos x, how would you find the derivative?

OpenStudy (mathmale):

Answer: the derivative of y = (tan x)/(cos x), following the quotient rule, is dy (cos x)*(sec x)^2 - (tan x)*(-sin x) -- = ----------------------------- dx (cos x)^2 If this is unclear, please ask questions, and then I (or someone else) will help you complete the work necessary to solve this homework problem.